How do you solve y=2x+2 in substation form
6x+y=4

Answer by mananth(16946)

y=2x+2 in substation form
6x+y=4
6x+y=4
substitute y
6x+2x+2= 4
8x+2=4
8x=2
x=2/8
x=1/4
plug x to get value of y
6*1/4 +y =4
y=4-3/2
y=5/2